At the beginning of 2020 you invest $3,000 in Microsoft (MSFT) stock and $5,000 in Proctor and Gamble (PG) stock.
Suppose you expect the monthly return to be 1.2% for MSTF and 0.5% for PG and the standard deviation of monthly returns to be 11.6% for MSTF and 6.5% for PG, and that the correlation of the returns on the two stock is 0.18.


- What are the expected monthly return on your portfolio and its stand dard deviation?

- Compare and discuss the return and standard deviation of the portfolio to the return and standard deviation of the two individual stocks.
Answer #1

Percentage Invested in Microsoft = (3000 / 8000) = 37.5%

Percentage Invested in Procter & gamble = (5000 / 8000) = 62.5%

Monthly return for Microsoft = 1.2%

Monthly return for Procter & gamble = 0.5%

Expected monthly return on portfolio = ( Weight of Stock A * Return of Stock A ) + ( Weight of Stock B * Return of Stock B )

= (0.375 * 1.2%) + (0.625 * 0.5%) = 0.76%

Expected standard deviation of portfolio = [ { ( Weight of Stock A * Std Deviation of Stock A) ^ 2 } + { ( Weight of Stock B * Std Deviation of Stock B) ^ 2 } + 2 * Weight of Stock A * Weight of Stock B * Correlation of Stock A & B * Std Deviation of Stock A * Std Deviation of Stock B ] ^ (1/2)

= 6.46%

Portfolio return will lie between Returns of 2 Stocks while Standard Deviation reduces to 6.46% on account of benefit of diversification